Two identical point charges, 10 cm apart in air, repel each other with a force of 1 newton. Find the magnitude of each charge.

1
Expert's answer
2022-03-16T18:30:04-0400

Force

"F=\frac{kq_1q_2}{r^2}"

"1=\frac{9\times 10^9 \times q_1q_2}{0.1^2}"

"q^2=\frac{0.1^2}{9\times10^9}"

"q=1.05\times10^{-6}C"

"q_1=q_2=q=1.05\times10^{-6}C"